OpenStudy (anonymous):

Given the equation y − 3 = 1/2(x + 6) in point-slope form, identify the equation of the same line in slope-intercept form.

OpenStudy (mertsj):

Use the distributive property on the right and then add 3 to both sides and you will have it in slope-intercept form which is y=mx+b

OpenStudy (anonymous):

what he said^ haha

OpenStudy (anonymous):

is it y= 1/2(x+6)

OpenStudy (anonymous):

hello....

OpenStudy (anonymous):

Distribute1/2 to x + 6

OpenStudy (anonymous):

i dont get it

OpenStudy (anonymous):

is my answer rite

OpenStudy (anonymous):

y= 1/2x+3

OpenStudy (mertsj):